Circular No. 6357

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996P IN NGC 5335 J. Mueller reports her discovery of a supernova (mag ∼ 17) found on

a IIIa-J plate taken on Mar. 25 UT by K. M. Rykoski and herself with the 1.2-m Oschin Schmidt telescope in the course of the second Palomar Sky Survey. SN 1996P is located 57′′.8 west and 22′′ south of the center of NGC 5335 (α = 13^{h}53^{m}.0, δ = +2^{o}50′, equinox 2000.0). No object appears at this position on original Sky Survey prints or on a second Sky Survey transparency taken by Mueller and D. J. Mendenhall on 1990 May 18.

B. Balick, A. Riera, Y. Terzian, and K. Xilouris report that an uncalibrated spectrogram (range 350–750 nm, resolution 0.3 nm, 2000-s exposure), obtained with the Hale 5-m telescope (+ double spectrograph), shows that emission lines of Hα, He II at 468.6 nm, and [O I] at 630.0 nm, as well as an absorption line of He I at 587.6 nm, appear to be present.

A. V. Filippenko and D. C. Leonard, University of California at Berkeley, report that inspection of a CCD spectrogram (range 430–700 nm, resolution 0.7 nm), obtained on Mar. 27 with the 3-m Shane reflector at Lick Observatory, reveals that SN 1996P is a supernova of type Ia, roughly 1.5–2 months past maximum brightness.

COMET C/1996 B2 (HYAKUTAKE) G.-P. Tozzi, Osservatorio Astrofisico di Arcetri (OAA), writes: “Pre-

liminary analysis of infrared observations of comet C/1996 B2 taken by F. Lisi and E. Oliva (OAA) and P. Patriarchi (CNR Gruppo Nazionale di Astronomia) at the TIRGO telescope on Mar. 25.93 UT shows a blob in the comet’s tail with a maximum intensity of ∼ 10 percent of the comet’s central condensation in the J band. The blob was at a projected distance of ∼ 18′′, corresponding to 1350 km at the comet distance. Later observations have shown that the blob moved 3′′.3 in the tail direction in ∼ 4 hr (corresponding projected linear distance 250 km; deduced velocity, ∼ 17 m/s, assuming that the blob is moving along the tail direction).”

COMET 22P/KOPFF Total visual magnitude estimates: Feb. 26.13 UT, 12.0 (K. Hornoch,

Lelekovice, Czech Republic, 0.35-m reflector); Mar. 17.45, 11.3 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 24.16, 11.7 (J. Carvajal, Teruel, Spain, 0.44-m reflector).

1996 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green